OpenStudy (anonymous):

Verify the identity. (1 − sin^2 x)(tan^2 x + 1) = 1 Use the Pythagorean and Reciprocal Identities, and simplify. (1-sin^2 x)(tan^2 x1) = (?) (sec^2 x) = (?) (1/cos^2 x) = 1

OpenStudy (anonymous):

x=0+2(pie)n

OpenStudy (mr.math):

\((1-\sin^2{x})(\tan^2{x}+1)=\cos^2{x}(\frac{1}{\cos^2{x}})=1\)

OpenStudy (anonymous):

What do I put in the space next to the sec?

OpenStudy (mr.math):

cos^2(x)